Calculate the MIRR for the following project. The discount rate is 10%.

t=0 t=1 t=2 t=3

-100 60 -30 150
Answer #1

FV of Cash inflows = [CFi * (1 + r)(n - i)]

= [60 * (1 + 0.10)(3 - 1)] + [150 * (1 + 0.10)(3 - 3)]

= 72.6 + 150 = 222.6

PV of Cash outflows = [CFi / (1 + r)n]

= 100 + [30 / (1 + 0.10)2]

= 100 + 24.79 = 124.79

MIRR = [FV of Cash inflows / PV of Cash outflows]1/n - 1

= [222.6 / 124.79]1/4 - 1

= [1.7837]0.25 - 1

= 1.1557 - 1 = 0.1557, or 15.57%